A firm in a perfectly competitive market has the following cost function:

C = 1/3q3 – 5q2

+ 30q +10

If the market-clearing price is 6, obtain the profit maximising level of output


1
Expert's answer
2021-04-06T07:11:15-0400

solution

cost function

c(q)="\frac{1}{3}q^3-5q^2+30q+10"

MCP=6


for maximum profit will be at the point where


P=MCP=MC


and

MC="\frac{dC(q)}{dq}"

"MC=\frac{d}{dq}(\frac{1}{3}q^3-5q^2+30q+10)"


MC="q^2-10q+30"


and MPC=6

so

"q^2-10q+30=6"


"q^2-10q+24=0"

by solving it

"q=4\space\space ,\space 6"


the profit maximizing level of output will be q=6